### v1

# Casimir-Lifshitz interaction between bodies integrated in a micro/nanoelectromechanical quantum damped oscillator

Abstract
A theory is proposed for the component of the Casimir-like force that arises between bodies embedded in a macroscopic quantum damped oscillator. When the oscillator’s parameters depend on the distance between the bodies, the oscillator-induced Casimir-like force is generally determined by a broad spectral range extending to high frequencies, limited by the frequency dispersion of the damping function. Here it is shown that there is a large class of systems in which the low-frequency range dominates the forces. This allows for the use of the Ohmic approximation, which is crucial for extending the theory to the lumped element description of fluctuation-induced forces in electrical circuits. Estimates of the circuit-induced Casimir-Lifshitz force suggest that under certain conditions it can be identified experimentally due to its dependence on various circuit elements.
Introduction.Casimir forces originate from quantum fluctuations of the electromagnetic vacuum in the presence of boundaries or other inhomogeneities [1, 2]. Extending the theory to both quantum and thermal electromagnetic fluctuations in inhomogeneous equilibrium condensed matter systems, led to a general theory of Casimir and van der Waals (or Casimir-Lifshitz) forces between atoms, molecules, and macroscopic bodies [3, 4, 5] (see also reviews [6, 7, 8, 9, 10, 11]).
Advances in micro- and nanophysics technologies, particularly in the development of mechanical transducers, atomic force microscopes, torsion pendulums, and oscillators, enabled detailed quantitative studies of Casimir-Lifshitz forces (see, for example, [12, 13, 14, 15, 10, 16, 17, 18, 19, 11, 20, 21, 22, 23, 24, 25, 26, 27] and references therein). These advances also drew significant attention to the influence of Casimir-Lifshitz forces on the performance of micro- and nanoelectromechanical devices, whose components often operate in close proximity [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49].
This paper presents a theory of fluctuation-induced interaction that arises between bodies embedded in quantum damped oscillators as encountered in micro- or nanoelectromechanical devices. The Casimir-like force considered here is driven by the oscillator through a parametric dependence of its frequency $\Omega$ and/or its damping function $\gamma(\omega)$ on the distance between the bodies. In electrical circuits, a similar Casimir-Lifshitz force component emerges from electromagnetic eigenmodes shaped by distance-dependent lumped elements. This component can, under certain conditions, be detected experimentally, typically as a correction to the main Casimir-Lifshitz force. It can be distinguished by its dependence on various circuit elements.
The position and momentum correlation functions, the partition function, and some related thermodynamic quantities of the quantum damped oscillator were studied in detail, mostly using the Zwanzig-Caldeira-Leggett model and assuming a bilinear coupling with the environment [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. These results can, in particular, be represented as sums or products over the Matsubara frequencies $\omega_{n}=2\pi nT/\hbar$.
The quantities in question can be divided into two groups, depending on the spectral range of fluctuations that dominate them. The first group, for example, includes the mean-square momentum fluctuation and the ground state energy. The corresponding sums diverge at high frequencies in the Ohmic approximation, because their convergence is ensured only by the frequency dispersion of the damping function $\gamma(i\omega_{n})$. This dispersion is determined by the spectral density of the environment [51, 53, 60], which decays rapidly with $\omega_{n}$ at high frequencies, thereby defining a characteristic frequency range $\omega_{n}\lesssim\omega_{c}$ that dominates the quantities in the first group.
By contrast, the quantities of the second group are primarily governed by fluctuations in the spectral region $\omega\lesssim\tilde{\Omega}$, where $\tilde{\Omega}\sim\Omega$ for small or moderate $\gamma$. When $\omega_{c}\gg\tilde{\Omega}$, that is applicable to many problems and is assumed below, the condition $\omega\lesssim\tilde{\Omega}$ is favorable for applying the Ohmic approximation or the lumped element approach. The mean squared position fluctuations of the damped oscillator and its specific heat are representative of this second group. Unlike the ground state energy, the specific heat of the quantum damped oscillator belongs to the second group since its behavior is determined by the temperature derivatives of the thermal part of the energy [58, 59, 61].
The oscillator-induced Casimir-like force provides an instructive example in this regard. Representing the distance derivative of the full free energy of the damped oscillator, the force belongs to the first group if the damping function depends on distance. In this case, the Casimir-like force includes a high-frequency contribution limited only by the damping function, whose frequency dispersion must be specified. The corresponding result is obtained below using the Drude model.
The role of frequency dependent parameters is crucial for extending the theory of oscillator-induced Casimir-like forces to electrical circuits with a lumped element description, when the eigenmodes have frequency-independent damping functions. The divergences that arise in this case present a major challenge to achieving this goal. This paper identifies possible solutions to the problem that occur under certain conditions.
Here we show that the low-frequency range dominates the forces only when the resonant frequency $\Omega$ depends on distance, while the damping function $\gamma$ does not. This makes a lumped element description of fluctuation-induced forces in electrical circuits possible. A similar description also applies to another situation, where either $\gamma$, or both $\Omega$ and $\gamma$, depend on distance, but only $\Omega$ depends on an additional parameter $\varkappa$. In this case, the force difference $f_{\Omega_{1}}-f_{\Omega_{2}}$, with $\Omega_{1,2}=\Omega(\varkappa_{1,2})$, is governed by the low-frequency range.
Oscillator-induced Casimir-like forces.Consider a quantum macroscopic damped oscillator, assuming that its resonant frequency $\Omega$ and the damping function $\gamma$ depend on the distance $d$ between the constituent interacting bodies. These quantities enter the quantum Langevin equation for the oscillator’s coordinate $M\ddot{\hat{Q}}+M\int_{0}^{t}d\tau\gamma(t-\tau)\dot{\hat{Q}}(\tau)+M\Omega^{2}\hat{Q}=\hat{\xi}(t)$.
It follows from the expression for the damped oscillator’s free energy $F(d)=-T\log Z(d)$, known in the Zwanzig-Caldeira-Leggett model together with the partition function $Z(d)$ [53, 55, 60, 63], that the interaction force $f=-\frac{\partial F}{\partial d}$ takes the form  [^1] See Supplemental Material for details of the calculations of the oscillator-induced Casimir-like force under various conditions, as well as of the Casimir-Lifshitz force induced by the RLC circuits in series and in parallel.
$f=f_{\Omega}+f_{\gamma}=-T\sideset{}{{}^{\prime}}{\sum}_{n=0}^{\infty}\frac{2\Omega(d)\frac{\partial\Omega(d)}{\partial d}+\omega_{n}\frac{\partial\gamma(i\omega_{n},d)}{\partial d}}{\omega_{n}^{2}+\gamma(i\omega_{n},d)\omega_{n}+\Omega^{2}(d)}.$ (1)
Here $\gamma(\omega)=\int_{0}^{\infty}\gamma(t)e^{i\omega t}dt$, the summation is taken over the Matsubara frequencies, and the prime at the summation sign indicates that the term with $n=0$ is taken with half weight.
In (1), each of the contributions to the interaction force, $f_{\Omega}$ or $f_{\gamma}$, describes repulsion when $\Omega$ or $\gamma(i\omega_{n})$, respectively, decreases with distance, and attraction when it increases.
The term $f_{\gamma}\propto\frac{\partial\gamma(i\omega_{n},d)}{\partial d}$ in (1) belongs to the first group because it is sensitive to and limited by the frequency dependence of $\gamma(i\omega_{n})$, and logarithmically diverges in the Ohmic approximation. By contrast, the term $f_{\Omega}\propto\frac{\partial\Omega(d)}{\partial d}$ in (1) is dominated by a comparatively low frequency range, which is $\omega_{n}\lesssim\Omega$ for small and moderate $\gamma$. Considering that $\omega_{c}\gg\Omega$, the force component $f_{\Omega}$ can be described in the Ohmic approximation and thus belongs to the second group.
Focusing on the oscillator- or circuit-induced forces shaped mostly by low-frequency fluctuations, we assume first that only $\Omega$ depends on $d$, while $\gamma(\omega)$ remains fixed. (Examples will be given below.) Then one obtains from (1) the following expression for the total fluctuation force $f(d)=f_{\Omega}(d)$ in the Ohmic approximation:
$f_{\Omega}(d)=-\left\{\dfrac{T}{\Omega}+\frac{i\hbar\Omega}{2\pi\sqrt{\Omega^{2}-\frac{\gamma^{2}}{4}}}\left[\psi\biggl{(}1+i\frac{\hbar\omega_{2}}{2\pi T}\biggr{)}\right.\right.\\ \left.\left.-\psi\left(1+i\frac{\hbar\omega_{1}}{2\pi T}\right)\right]\right\}\dfrac{\partial\Omega}{\partial d},$ (2)
where the quantities $i\omega_{1,2}=\frac{\gamma}{2}\pm i\sqrt{\Omega^{2}-\frac{\gamma^{2}}{4}}$ enter the arguments of the digamma functions $\psi(x)$.
For weak dissipation $\gamma^{2}\ll 4\Omega^{2}$, the force (2) is given by the derivative of the free energy of the quantum harmonic oscillator $f=-\frac{1}{2}\coth\frac{\hbar\Omega}{2T}\frac{\partial\Omega}{\partial d}$. At low temperatures $T\ll\hbar\Omega$, this reduces to the contribution from the zero-point energy $f_{\Omega}=-\frac{\hbar}{2}\frac{\partial\Omega}{\partial d}$. At high temperatures $\hbar\Omega\ll T$, the first term in (2) dominates the force.
If $\gamma$ is not negligibly small, the dominating term that follows from (2) in the low temperature limit is
$f_{\Omega}=-\frac{\hbar\Omega\frac{d\Omega}{dd}}{\pi\sqrt{\gamma^{2}-4\Omega^{2}}}\ln\frac{\gamma+\sqrt{\gamma^{2}-4\Omega^{2}}}{\gamma-\sqrt{\gamma^{2}-4\Omega^{2}}}.$ (3)
In the limit of a strong dissipation, $4\Omega^{2}\ll\gamma^{2}$, the frequency $i\omega_{2}\approx\frac{\Omega^{2}}{\gamma}$ becomes anomalously small with increasing $\gamma$, while another one linearly increases $i\omega_{1}\approx\gamma$. The low temperature range is $T\ll\frac{\Omega^{2}}{\gamma}$ in this case, while the value of $\gamma$ is confined by the condition $\gamma\ll\omega_{c}$.
Figure 1: Simplest RLC circuits connected in series (a) and in parallel (b).
If $\gamma(\omega,d)$ depends on distance $d$, possibly together with $\Omega$, then an additional term $f_{\gamma}\propto\frac{\partial\gamma}{\partial d}$ appears in the expression of the total fluctuation force $f=f_{\Omega}+f_{\gamma}$, and a wide spectral range that includes the high frequences $\omega\lesssim\omega_{c}$ forms $f_{\gamma}$. Let $\Omega$, however, be dependent on an additional parameter $\varkappa$, while $\gamma(i\omega_{n},d)$ is not. One can see from (1) that the difference $f_{\gamma}(d,\varkappa_{1})-f_{\gamma}(d,\varkappa_{2})$ belongs to the second group, whereas the individual quantities $f_{\gamma}(d,\varkappa_{1,2})$ enter the first group.
Introducing the quantity $\tilde{f}(d,\varkappa)$, that satisfies the relation $\tilde{f}(d,\varkappa_{1})-\tilde{f}(d,\varkappa_{2})=f(d,\varkappa_{1})-f(d,\varkappa_{2})$ and retains only those terms from $f(d,\varkappa)$, which generally do not cancel out in this difference, we obtain within the Ohmic approximation
[…]
For describing the total oscillator-induced Casimir-like force in the presence of distance-dependent $\gamma$, the frequency dependence of $\gamma$ must be specified in (1). Consider here the Drude-like model $\gamma(\omega)=\frac{\gamma_{0}\omega_{D}}{\omega_{D}-i\omega}$, which is commonly used in studying the damped oscillator properties [52, 53, 55], and assume that generally all three model parameters, $\Omega$, $\gamma_{0}$ and $\omega_{D}$, depend on $d$. A pronounced frequency dependence of $\gamma(i\omega,d)$ shows up in the model in the range $\omega\gtrsim\omega_{D}$, so that $\omega_{D}$ plays in (1) the role of an effective cutoff frequency $\omega_{c}$.
In the Drude model each term involving $\frac{\partial\gamma_{0}}{\partial d}$ or $\frac{\partial\omega_{D}}{\partial d}$ describes repulsion when the parameter $\gamma_{0}$ or $\omega_{D}$ decreases with $d$, and attraction when it increases.
When the frequency dispersion of $\gamma$ is disregarded, the damped oscillator has two complex eigenfrequencies $\omega_{1,2}$, as used in (2) and (4), whereas in the Drude model it has three, $\omega_{1,2,3}$. If $\omega_{D}\gg\Omega,\gamma_{0}$, as is assumed here, these frequencies can be readily obtained using an expansion in powers of the small parameters $\frac{\Omega}{\omega_{D}}$ and $\frac{\gamma_{0}}{\omega_{D}}$ [52]. To first order, $i\omega_{1,2}=\frac{\gamma_{0}}{2}\pm i\sqrt{\Omega^{2}-\frac{\gamma_{0}^{2}}{4}}$ and $i\omega_{3}=\omega_{D}-\gamma_{0}\gg|\omega_{1,2}|$. In this case, the total fluctuation force is $f(d)=\tilde{f}(d)+\Delta f_{\gamma}(d)$ [64]. The first term is defined in (4), where $\gamma_{0}$ must be substituted for $\gamma$, and the second term contains the third eigenfrequency and the parameter $\omega_{D}$:
$\Delta f_{\gamma}(d)=-\dfrac{\hbar}{2\pi}\psi\left(1+\frac{\hbar(\omega_{D}-\gamma_{0})}{2\pi T}\right)\dfrac{\partial\gamma_{0}}{\partial d}+\\ \dfrac{\hbar}{2\pi}\left[\psi\left(1+\frac{\hbar(\omega_{D}-\gamma_{0})}{2\pi T}\right)-\psi\left(1+\frac{\hbar\omega_{D}}{2\pi T}\right)\right]\dfrac{\partial\omega_{D}}{\partial d}.$ (5)
In the high temperature range $\omega_{D}\gg T\gg\Omega,\gamma_{0}$ one finds from (4) and (5)
$f=-\dfrac{T}{\Omega}\dfrac{\partial\Omega}{\partial d}-\dfrac{\hbar}{2\pi}\dfrac{\partial\gamma_{0}}{\partial d}\ln\frac{\hbar\omega_{D}}{2\pi T}-\dfrac{\hbar\gamma_{0}}{2\pi\omega_{D}}\dfrac{\partial\omega_{D}}{\partial d}.$ (6)
Three contributions to the interaction force in (6) arise from the distance dependence of $\Omega$, $\gamma_{0}$ and $\omega_{D}$ and can generally have the opposite signs. The large argument under the logarithm sign in (6) reflects the logarithmic divergence of the sum in (1), which appears when the frequency dispersion of $\gamma$ is neglected.
Casimir-Lifshitz forces in circuits.Assuming that dissipation in electrical circuits is described within the Zwanzig-Caldeira-Leggett model [65, 66, 67], we apply the above results to the two simplest RLC circuits: one in series and the other in parallel, as depicted in Fig. 1.
In the RLC circuit connected in series, the charge $Q$ can be taken as the system’s generalized coordinate [68]. The corresponding canonical momentum is the current-induced magnetic flux $\Phi=L\dot{Q}$. The equation of motion for the circuit, $L\ddot{Q}+R\dot{Q}+\frac{Q}{C}={\cal E}_{ext}(t)$, with the external voltage ${\cal E}_{ext}(t)$, has the same form as that of the damped harmonic oscillator. In this analogy, $M\rightarrow L$, $\gamma\rightarrow\frac{R}{L}$ and $\Omega\rightarrow\frac{1}{\sqrt{LC}}$. Note that $\gamma$ here represents the frequency-independent Ohmic damping.
If the capacitance $C$ depends on the distance $d$, while $R$ and $L$ do not, then only the oscillator frequency $\Omega$ depends on $d$, whereas $\gamma$ and $M$ do not. Therefore, the contribution $f_{RLC}$ of electromagnetic fluctuations in the series circuit to the Casimir-Lifshitz force is determined by equation (2), with parameters $\Omega_{LC}=\frac{1}{\sqrt{LC}}$ and $\gamma=\frac{R}{L}$. Alternatively, if $C$ is the only circuit parameter that depends on a quantity $\varkappa$, while some other parameters (possibly including $C$) depend on $d$, then the finite force difference is $\tilde{f}_{RLC_{1}}-\tilde{f}_{RLC_{2}}$, where $\tilde{f}_{RLC}$ is defined in (4) with $\Omega_{LC_{1,2}}=\frac{1}{\sqrt{LC_{1,2}}}$, $\gamma=\frac{R}{L}$ and $C_{1,2}=C(\varkappa_{1,2})$.
In the RLC circuit connected in parallel, the generalized coordinate is $\Phi=\int_{-\infty}^{t}V(t^{\prime})dt^{\prime}$, where $\Phi$ is the magnetic flux. The capacitor charge $Q_{C}=C\dot{\Phi}=CV$ serves as the canonical momentum in this case [65, 67, 69]. The corresponding equation of motion, $C\ddot{\Phi}(t)+\dfrac{1}{R}\dot{\Phi}(t)+\dfrac{1}{L}\Phi(t)=J_{\text{ext}}(t)$, coincides with that for a damped harmonic oscillator, where $M\rightarrow C$, $\gamma\rightarrow\frac{1}{RC}$ and $\Omega\rightarrow\frac{1}{\sqrt{LC}}$. Here, as before, $\gamma$ is frequency-independent.
If the inductance $L$ depends on distance $d$, while $R$ and $C$ do not, then only the oscillator frequency $\Omega$ in the parallel RLC circuit depends on $d$, whereas $\gamma$ and $M$ remain unchanged. Therefore, the contribution $f_{RLC}$ of the circuit to the Casimir-Lifshitz force in the parallel configuration is determined by (2), with parameters $\Omega_{LC}=\tfrac{1}{\sqrt{LC}}$ and $\gamma=\tfrac{1}{RC}$. Alternatively, if $L$ is the only circuit parameter that depends on a quantity $\varkappa$, while a few circuit parameters (possibly, including $L$) depend on $d$, then the force difference is $f_{RL_{1}C}-f_{RL_{2}C}=\tilde{f}_{RL_{1}C}-\tilde{f}_{RL_{2}C}$, where $\tilde{f}_{RLC}$ is defined in (4), $\Omega_{LC_{1,2}}=\tfrac{1}{\sqrt{L_{1,2}C}}$, $\gamma=\tfrac{1}{RC}$ and $L_{1,2}=L(\varkappa_{1,2})$.
We estimate the Casimir-Lifshitz force component $f_{RLC}$ for the RLC loop in series involving the distance-dependent capacitance, and examine further the ratio $r=\frac{f_{RLC}}{f_{Cas}}$, which compares this force component to the main Casimir-Lifshitz force $f_{Cas}$ between the same bodies under similar conditions. Consider first the planar capacitor, for which the resonant frequency of the circuit is $\Omega^{pc}_{LC}=\sqrt{\frac{d}{\varepsilon_{0}\varepsilon LS}}$, where $S$ and $d$ denote the contact area of the conducting plates and the distance between them, respectively; $\varepsilon_{0}$ is the vacuum permittivity. Focusing on the Casimir problem [1], we put $\varepsilon=1$ for the dielectric permittivity of the interlayer between the plates.
The force at high temperatures described by the first term in (2), $f_{T}=-\frac{T}{\Omega}\frac{\partial\Omega}{\partial d}$, takes the form $f^{pc}_{RLC,T}=-\frac{T}{2d}$ in this case. At low temperatures, one obtains from (3) for the interaction force at weak ($R^{2}\ll\frac{4Ld}{\varepsilon S}$) and strong ($R^{2}\gg\frac{4Ld}{\varepsilon S}$) dissipation, respectively:
$f^{pc}_{RLC,0}=-\frac{\hbar}{4\sqrt{\varepsilon_{0}LSd}}+\frac{\hbar R}{4\pi Ld},\quad T\ll\frac{\hbar\sqrt{d}}{2\pi\sqrt{\varepsilon_{0}SL}},$ (7)
$f^{pc}_{RLC,0}=-\frac{\hbar}{2\pi\varepsilon_{0}SR}\log\frac{\varepsilon_{0}SR^{2}}{Ld},\quad T\ll\frac{\hbar d}{2\pi\varepsilon_{0}SR}.$ (8)
Given the applicability conditions of the lumped element approach, the strength of dissipation in (8) is confined by the requirement $i\omega_{1}\approx\frac{R}{L}\ll\min(\omega_{c},c/r_{0})$, where $r_{0}$ is the typical size of a circuit element.
Furthermore, the conventional Casimir force, $f_{Cas}=-\frac{\pi^{2}\hbar cS}{240d^{4}}$, is the main fluctuation force component between metal plates in the strong retardation regime at low temperatures. At high temperatures, $T\gg\hbar c/d$, the force becomes $f_{Cas,T}=-\frac{\zeta(3)TS}{8\pi d^{3}}$. This expression follows from the Lifshitz theory with the Drude model  [^2] As was noted long ago (see p.263 in Ref. [86]), this is half of what follows from the plasma model, since the latter fails to account for the complete suppression of the TM polarization contribution at high temperatures. This fact was later rediscovered and drew considerable attention due to a reported discrepancy between precision measurements and predictions from Lifshitz theory [90, 91, 77, 26, 92].
One finds from here the relative weights $r^{pc}_{0}$ and $r^{pc}_{T}$ of the circuit-induced Casimir-Lifshitz force between the plates of the planar capacitor in the dissipationless limit, $\hbar\gamma\ll\hbar\Omega,T$, at low and high temperatures:
$r_{0}^{pc}=\frac{60d^{7/2}}{\pi^{2}c\sqrt{\varepsilon_{0}LS^{3}}}=\frac{60}{\pi^{2}}\left(\frac{\Omega^{pc}_{LC}(d)}{c/d}\right)\frac{d^{2}}{S},$ (9)
$r_{\phantom{}{}_{T}}^{pc}=\frac{4\pi d^{2}}{\zeta(3)S}.$ (10)
As seen in (9) and (10), the relative weight involves two small parameters at low temperature range and only one at high. The quantities $r^{pc}_{0}$ and $r_{\phantom{}{}_{T}}^{pc}$, are both governed by the small geometric factor $\frac{d^{2}}{S}$. It reflects the different dependence of the forces on the separation between the bodies and their characteristic sizes. Its smallness justifies neglecting edge effects at the boundaries of the plates.
Another small parameter, $\frac{\Omega^{pc}_{LC}(d)}{c/d}$, that appears in $r_{0}^{pc}$ is the ratio of the oscillator frequency $\Omega^{pc}_{LC}$ and the characteristic frequency $c/d$ of quantum fluctuations forming the Casimir force. This ratio is typically small throughout the entire applicability domain of the Casimir result $f_{Cas}$. After switching over from low to high temperatures, both characteristic frequencies, $\Omega^{pc}_{LC}$ and $d/c$, present in (9) are effectively replaced by $T$, which cancels out the frequencies in the relative weight, leading to (10), up to a numerical factor.
As the ratio $\frac{d^{2}}{S}$ decreases from $0.04$ to $2.5\!\cdot\!10^{-3}$, the relative weight $r^{pc}_{\phantom{}{}_{T}}$ in (10) drops from $0.42$ to $0.03$. The corresponding measurement accuracy requirements indicate that the force component under consideration can generally be detected at high temperatures.
For estimating the relative weights for a metal sphere and a flat metal plate often used in experimental setups [12, 13, 71, 30, 14, 14, 72, 73, 74, 75, 76, 18, 77, 22, 41, 23, 25, 47], the capacitance could be calculated numerically [78, 79, 80, 81]. However, we find it sufficient for our purposes to rely on a simple interpolating expression, $C^{sp-p}(d,R_{sp})=4\pi\varepsilon_{0}R_{sp}\left[1+\frac{1}{2}\log\left(1+\frac{R_{sp}}{d}\right)\right]$, which provides a good approximation for $d\lesssim R_{sp}$ [80]. Here, $R_{sp}$ is the radius of the sphere, and $d$ is the minimum sphere-plate distance. In the dissipationless limit and at low temperatures, $T\ll\hbar\Omega^{sp-p}_{LC}$, the circuit-induced fluctuation force between the sphere and the plate, associated with the corresponding $\Omega^{sp-p}_{LC}=\frac{1}{\sqrt{LC^{sp-p}}}$, is
$f^{sp-p}_{LC}=-\frac{\hbar\Omega^{sp-p}_{LC}(d)}{8d\bigl{(}1+\frac{d}{R_{sp}}\bigr{)}\bigl{[}1+0.5\log\bigl{(}1+\frac{R_{sp}}{d}\bigr{)}\bigr{]}},$ (11)
while at high temperatures $T\gg\hbar\Omega^{sp-p}_{LC}$ one obtains
$f^{sp-p}_{LC}=-\frac{T}{4d\bigl{(}1+\frac{d}{R_{sp}}\bigr{)}\bigl{[}1+0.5\log(1+\frac{R_{sp}}{d})\bigr{]}}.$ (12)
The main Casimir-Lifshitz attraction between the sphere and the plate, in the proximity force approximation, is given by $f^{sp-p}=-\pi^{3}\hbar cR_{sp}/360d^{3}$ at $T\ll\hbar c/d$, and $f^{sp-p}=-\zeta(3)TR_{sp}/8d^{2}$ at $T\gg\hbar c/d$. Known refinements of these simple expressions [82, 26] do not affect the subsequent conclusions.
Thus, in the dissipationless limit, the relative weights of the circuit-induced fluctuation force between the metallic sphere and plate at low ($T\ll\hbar c/d$) and high ($T\gg\hbar c/d$) temperatures are, respectively
$r^{sp-p}_{0}=\left(\frac{\Omega^{sp-p}_{LC}}{c/d}\right)\frac{1.45}{\bigl{(}\frac{R_{sp}}{d}+1\bigr{)}\bigl{[}1+0.5\log\bigl{(}\frac{R_{sp}}{d}+1\bigr{)}\bigr{]}},$ (13)
$r^{sp-p}_{T}=\frac{1.66}{\bigl{(}\frac{R_{sp}}{d}+1\bigr{)}\bigl{[}1+0.5\log\bigl{(}\frac{R_{sp}}{d}+1\bigr{)}\bigr{]}}.$ (14)
The factor $\Omega^{sp-p}_{LC}/(c/d)$ that enters (13) is typically small within the applicability domain of the Casimir result. However, this factor is canceled out in (14). The geometric parameter $\frac{d}{R_{sp}}$ can vary here within comparatively wide limits. The relative weight $r^{sp-p}_{T}$ decreases from $0.5$ to $0.02$ as $\frac{d}{R_{sp}}$ decreases from $0.75$ to $0.035$. This shows a possibility to detect the circuit-induced force $f^{sp-p}_{LC}$ under optimal conditions. The presence of multiple distance-dependent eigenmodes could enhance the effect.
Discussion and conclusions.Components of the Casimir-Lifshitz forces induced in electrical circuits were previously examined in Ref. [83]. That paper addresses a number of important aspects of the issue under discussion and presents a series of valuable and useful findings. However, the main results and statements in Ref. [83] concerning the circuit-induced component of the Casimir-Lifshitz force are in contradiction to those obtained here. This relates both to the formulas for the forces in the circuits and to the more general results and conclusions.
One such conclusion in Ref. [83] is that, in the dissipationless limit at zero temperature, the interaction potential differs from the zero-point energy. The other conclusion is that the divergence of the interaction potential, as described within the lumped element approach, also entails the divergence of the interaction force, even when it is only the resonant frequency that depends on distance. Specifically, the expression for interaction force in the series RLC loop, as presented in Ref. [83], is divergent in the case of a distance-dependent $C$ at constant $R$ and $L$. This contradicts our fomulas (7) and (8), as well as the more general results (2) and (3).
The main reason for the contradictions is that the force component between the capacitor plates, as considered in Ref. [83], is attributed solely to charge fluctuations on the plates themselves, without taking into account other contributions. However, this only represents part of the total Casimir-Lifshitz force induced by the circuit, since changing the plate separation alters the impedance of the entire circuit. Consequently, the circuit’s free energy dependence on distance is determined by fluctuations in all lumped circuit elements. In this regard, it is worth recalling that the fluctuation-induced interaction between oscillators or atoms originates from both potential and kinetic fluctuational energy terms [84, 85].
Before the technological breakthrough in the field, the fluctuation-induced forces in the series RLC circuit were considered a purely methodological issue [86],  [^3] See also the footnote on p.407 in Ref. [8]., primarily in connection with Ref. [88]. While the fluctuation-induced free energy in the series RLC circuit diverges, it was shown that the difference between free energies, taken at different capacitance values, is finite. A finite additional component of the fluctuation force that was identified in Ref. [86] for the circuit, coincides with (7), (8) up to a minor typo  [^4] In the limit of large dissipation, described by equation (5.422) on page 248 in Ref. [86], the resistance $R$ under the logarithm sign must be squared.. Those early results were obtained by simply expressing the fluctuation energy in the RLC circuit as the average of $\frac{L\dot{Q}^{2}}{2}+\frac{Q^{2}}{2C}$.
This paper proposes a microscopic theory of Casimir-like forces induced by the quantum damped oscillator. The application of this approach to bodies integrated in micro/nanoelectromechanical systems reveals the conditions for a lumped description of the circuit-induced Casimir-Lifshitz forces. While the free energy diverges under those conditions, the circuit-induced component of the force has been identified and shown to be finite.
The results obtained can be straightforwardly extended to other problems involving damped oscillators or electrical circuits, where a derivative of free energy $F$ with respect to a parameter other than the interbody distance is of interest. The only requirement is that the dependence of $F$ on this parameter is entirely determined by the dependence of $\Omega$ and $\gamma$ on it. For example, an interaction torque arises when $\Omega$ and/or $\gamma$ depend on the misorientation angle $\theta$ between the interacting bodies. This torque can be derived from (1)–(6) by replacing $\frac{\partial\Omega}{\partial d}$ and $\frac{\partial\gamma}{\partial d}$ with $\frac{\partial\Omega}{\partial\theta}$ and $\frac{\partial\gamma}{\partial\theta}$, respectively.
The work has been carried out within the state task of ISSP RAS.
Supplemental Material
Yu. S. Barash
Osipyan Institute of Solid State Physics RAS, Chernogolovka, Moscow District, 2 Academician Ossipyan Street, 142432 Russia
S1 Casimir-like forces induced by the damped oscillator
The Zwanzig-Caldeira-Leggett model describes a harmonic oscillator coupled to a thermal bath, represented by a large (or infinite) set of other harmonic oscillators. In the absence of external forces, the Hamiltonian of the full system involving bilinear coupling with the environment takes the form
$\hat{H}(d)=\dfrac{\hat{P}^{2}}{2M(d)}+\dfrac{1}{2}M(d)\Omega^{2}(d)\hat{Q}^{2}+\\ \sum_{\alpha}\left[\dfrac{\hat{p}_{\alpha}^{2}}{2m_{\alpha}}+\dfrac{1}{2}m_{\alpha}\omega_{\alpha}^{2}\left(\hat{q}_{\alpha}-\dfrac{C_{\alpha}(d)}{m_{\alpha}\omega_{\alpha}^{2}}\hat{Q}\right)^{2}\right].$ (S1)
Here, we additionally assume that the resonant frequency $\Omega(d)$ of the central oscillator, as well as, generally speaking, its mass $M(d)$ and its coupling constants with the surrounding oscillators $C_{\alpha}(d)$, depend on the distance $d$ between the bodies embedded in the scillator. This type of dependence can occur, in particular, in damped oscillators of micro-/nanoelectromechanical electrical circuits.
The interaction force is calculated in a statistically equilibrium state based on the Hamiltonian (S1) of the complete closed system, including the environment. The damping function enters the expression for the force only after statistical averaging. Since a small change in the distance between bodies, $\delta d$, is assumed to occur at nearly zero speed and at constant temperature, all changes in free energy, $\delta F$, should be attributed to the work of the interaction force along the path $\delta d$. After averaging, there is no dissipation in equilibrium.
Under these conditions, the interaction force is given by $f(d)=-\left\langle\frac{\partial\hat{H}(d)}{\partial d}\right\rangle_{T}$, which, according to general results of statistical physics, is equivalent to $f=-\left(\frac{\partial F}{\partial d}\right)_{T}$. This can also be verified by a direct calculation, which straightforwardly generalizes the calculation of the free energy of a damped oscillator in Ref. [63] performed by integrating over the interaction constant.
When performing statistical averaging of the derivative of the Hamiltonian with respect to distance, the resulting spectral densities of the symmetrized correlation functions $\left(P^{2}\right)_{\omega}$, $\left(Q^{2}\right)_{\omega}$, and $\bigl{(}q_{\alpha}Q\bigr{)}_{\omega}$ can be described using the following fluctuation-dissipation relations
$\bigl{(}Q^{2}\bigr{)}_{\omega}=\hbar\coth\left(\dfrac{1}{2}\beta\hbar\omega\right)\operatorname{\mathrm{Im}}\Bigl{[}\chi(\omega)\Bigr{]},$ (S2)
$\bigl{(}q_{\alpha}Q\bigr{)}_{\omega}=\hbar\coth\left(\dfrac{1}{2}\beta\hbar\omega\right)C_{\alpha}\operatorname{\mathrm{Im}}\Bigl{[}\chi(\omega)\chi_{\alpha}(\omega)\Bigr{]},$ (S3)
$\bigl{(}P^{2}\bigr{)}_{\omega}=\hbar\coth\left(\dfrac{1}{2}\beta\hbar\omega\right)M^{2}\omega^{2}\operatorname{\mathrm{Im}}\Bigl{[}\chi(\omega)\Bigr{]}.$ (S4)
Equations (S2)-(S4) involve the linear susceptibility of the damped oscillator
$\chi(\omega,d)=\dfrac{1}{M(d)\left[\left(\Omega^{2}(d)-\omega^{2}\right)-i\omega\gamma(\omega,d)\right]},$ (S5)
which depends on $d$ through the distance dependence of the damping function $\gamma(\omega,d)$, the resonant frequency $\Omega(d)$, and the mass $M(d)$.
The damping function
$\gamma(\omega,d)=-\,\dfrac{i}{\omega M(d)}\sum_{\alpha}C_{\alpha}^{2}(d)\left[\chi_{\alpha}(\omega)-\chi_{\alpha}(0)\right]$ (S6)
depends on the interbody distance when the oscillator mass and/or its coupling constants with the thermal bath modes exhibit such a dependence.
At the same time, the susceptibility of the original individual reservoir mode (i.e., of the corresponding free oscillator) that enters (S6) and (S3) does not depend on $d$:
$\chi_{\alpha}(\omega)=\dfrac{1}{m_{\alpha}}\dfrac{1}{\omega_{\alpha}^{2}-\omega^{2}-i\omega\varepsilon},\quad\varepsilon\to+0\,.$ (S7)
The calculation schematically outlined above leads to the equality $f=-\left(\frac{\partial F}{\partial d}\right)_{T}$, which involves the following expression for the free energy $F=-T\log Z$ of a damped oscillator
$F=T\ln\left[\dfrac{\hbar\Omega}{T}\prod\limits_{n=1}^{\infty}\left(1+\dfrac{\Omega^{2}}{\omega_{n}^{2}}+\dfrac{\gamma(i\omega_{n})}{\omega_{n}}\right)\right].$ (S8)
Free energy (S8) is known together with the corresponding partition function [53, 55, 60, 63]. As a result, we arrive at the interaction force (1), given in the main text of this paper.
S2 Casimir-like forces in the Ohmic approximation
Suppose that the resonant frequency of the oscillator $\Omega$ varies with a certain parameter $\varkappa$, while the damping function $\gamma(\omega)$ does not depend on $\varkappa$. Let’s consider the difference between free energies $F_{1,2}=F(\varkappa_{1,2})$, corresponding to two values $\varkappa_{1,2}$. It follows from (S8)
$F_{2}-F_{1}=T\ln\left[\dfrac{\Omega_{2}}{\Omega_{1}}\prod\limits_{n=1}^{\infty}\dfrac{\omega_{n}^{2}+\omega_{n}\gamma(i\omega_{n})+\Omega_{2}^{2}}{\omega_{n}^{2}+\omega_{n}\gamma(i\omega_{n})+\Omega_{1}^{2}}\right],$ (S9)
where $\Omega_{1,2}=\Omega(\varkappa_{1,2})$.
Unlike (S8), the expression on the right-hand side of (S9) converges within the Ohmic approximation. This becomes obvious when the relation (S9) is rewritten in the form
$F_{2}-F_{1}=T\sideset{}{{}^{\prime}}{\sum}_{n=0}^{\infty}\ln\left(1+\dfrac{\Omega_{2}^{2}-\Omega_{1}^{2}}{\omega_{n}^{2}+\omega_{n}\gamma(i\omega_{n})+\Omega_{1}^{2}}\right).$ (S10)
Here, the prime at the summation sign indicates that the term with $n=0$ is taken with half-weight.
Within the Ohmic approximation adopted in this section, we use the roots $\omega_{1,2}$ of the equation
$\omega^{2}+i\gamma\omega-\Omega^{2}=0,$ (S11)
which represent the oscillator’s complex eigenfrequencies in this approach.
One gets
$i\omega_{1,2}=\frac{\gamma}{2}\pm i\sqrt{\Omega^{2}-\frac{\gamma^{2}}{4}}.$ (S12)
Given equation (S12) and the Matsubara frequency $\omega_{n}=\tfrac{2\pi T}{\hbar}n$, the difference in free energies (S9) can be written as
$F_{2}-F_{1}=T\ln\left[\dfrac{\Omega_{2}}{\Omega_{1}}\prod\limits_{n=1}^{\infty}\dfrac{\left(n-\frac{\hbar\omega_{1}(\Omega_{2})}{2\pi T}\right)\left(n-\frac{\hbar\omega_{2}(\Omega_{2})}{2\pi T}\right)}{\left(n-\frac{\hbar\omega_{1}(\Omega_{1})}{2\pi T}\right)\left(n-\frac{\hbar\omega_{2}(\Omega_{1})}{2\pi T}\right)}\right].$ (S13)
Equality $\omega_{1}(\Omega_{2})+\omega_{2}(\Omega_{2})=\omega_{1}(\Omega_{1})+\omega_{2}(\Omega_{1})$ ensures the convergence of the infinite product in (S13), allowing us to apply the summation formula (1.3.8) on p. 7 of Ref. [93]. As a result, the quantity $F_{2}-F_{1}$ in (S13) is expressed in terms of Gamma functions:
$F_{2}-F_{1}=T\ln\left[\dfrac{\Omega_{2}}{\Omega_{1}}\dfrac{\Gamma(1-\frac{\hbar\omega_{1}(\Omega_{1})}{2\pi T})\Gamma(1-\frac{\hbar\omega_{2}(\Omega_{1})}{2\pi T})}{\Gamma(1-\frac{\hbar\omega_{1}(\Omega_{2})}{2\pi T})\Gamma(1-\frac{\hbar\omega_{2}(\Omega_{2})}{2\pi T})}\right].$ (S14)
When moving from free energy to interaction force, two cases should be distinguished, both permitting the use of the Ohmic approximation. In the first case, the parameters $\varkappa$ and $d$ vary independently. In the second case, they coincide, $\varkappa\equiv d$.
Let only the frequency $\Omega$ depend on $\varkappa$ in the first case, while the damping function $\gamma$ depends on the distance $d$, possibly together with $\Omega$. Then the relatively low frequency range makes the dominating contribution to the difference of the interaction forces $f(\varkappa_{2},d)-f(\varkappa_{1},d)=-\frac{\partial}{\partial d}(F_{2}-F_{1})$. From (S14) and (S12), it follows that this difference is described by expression (4) of the main text. The quantity $\tilde{f}(d,\kappa)$ in (4) satisfies the relation $\tilde{f}(d,\varkappa_{2})-\tilde{f}(d,\varkappa_{1})=f(d,\varkappa_{2})-f(d,\varkappa_{1})$ and includes only those terms from $f(d,\kappa)$, which do not, in general, cancel out in the difference.
Considering the second case $d=\varkappa$, one assumes that only the resonant frequency $\Omega$ depends on $d$. Since not only the finite difference (S14) between free energies, but also the derivative of the free energy with respect to $\varkappa$ can be described within the Ohmic approximation, one may set $d_{2}=d_{1}+\delta d$ in (S13) and (S14) and arrive at expression (2) of the main text for the total Casimir-like force, $f=-\left(\tfrac{\partial F}{\partial d}\right)_{T}$, induced by the oscillator.
In expanded form, this expression reads:
$f=-\dfrac{T}{\Omega}\dfrac{\partial\Omega}{\partial d}-\left[\psi\left(1+\frac{\hbar\gamma}{4\pi T}-\frac{i\hbar}{2\pi T}\sqrt{\Omega^{2}-\frac{\gamma^{2}}{4}}\right)-\psi\left(1+\frac{\hbar\gamma}{4\pi T}+\frac{i\hbar}{2\pi T}\sqrt{\Omega^{2}-\frac{\gamma^{2}}{4}}\right)\right]\dfrac{i\hbar\Omega\frac{\partial\Omega}{\partial d}}{2\pi\sqrt{\Omega^{2}-\frac{\gamma^{2}}{4}}}.$ (S15)
In the limit of weak dissipation, $\gamma\ll\Omega,T$, we obtain from (S15) keeping only the first-order terms in the expansion in powers of $\gamma$ and using the relation [94], $\operatorname{\mathrm{Im}}\psi(1+iy)=-\tfrac{1}{2y}+\tfrac{\pi}{2}\coth(\pi y)$:
$f=-\left[\frac{\hbar}{2}\coth\frac{\hbar\Omega}{2T}+\frac{\hbar^{2}\gamma}{4\pi^{2}T}\operatorname{\mathrm{Im}}\left(\psi^{(1)}\biggl{(}1+\dfrac{i\hbar\Omega}{2\pi T}\biggr{)}\right)\right]\frac{\partial\Omega}{\partial d},$ (S16)
where $\psi^{(1)}(z)=\tfrac{d\psi(z)}{dz}$.
In the high-temperature limit, we expand the $\psi$-functions in (S15) in powers of small parameters, using the relations [94] $\psi(1+z)=-\gamma_{eil}+\tfrac{\pi^{2}}{6}z+\ldots\,$ and $\psi^{(1)}(1+z)=\tfrac{\pi^{2}}{6}-2\zeta(3)z+\ldots$, which are valid at small z. Here $\gamma_{eil}=0.577\ldots$ is the Euler’s constant. This leads to the equality
$f=-\left(\dfrac{T}{\Omega}+\dfrac{\hbar^{2}\Omega}{12T}\right)\dfrac{\partial\Omega}{\partial d}.$ (S17)
In the low temperature limit, it is convenient to transform (S15) using the relation $\psi(1+z)=\tfrac{1}{z}+\psi(z)$ and to consider further the first terms of the asymptotic expansion $\psi(z)=\ln z-\tfrac{1}{2z}+\ldots$. This results in
$f=-\frac{\hbar\Omega\frac{d\Omega}{dd}}{2\pi\sqrt{\frac{\gamma^{2}}{4}-\Omega^{2}}}\ln\frac{\frac{\gamma}{2}+\sqrt{\frac{\gamma^{2}}{4}-\Omega^{2}}}{\frac{\gamma}{2}-\sqrt{\frac{\gamma^{2}}{4}-\Omega^{2}}}.$ (S18)
When $2\Omega\geq\gamma$, it is convenient to transform (S18) into
$f=-\frac{\hbar\Omega\frac{d\Omega}{dd}}{2\sqrt{\Omega^{2}-\frac{\gamma^{2}}{4}}}\left(1-\frac{2}{\pi}\arctan\frac{\frac{\gamma}{2\Omega}}{\sqrt{1-\frac{\gamma^{2}}{4\Omega^{2}}}}\right)$ (S19)
S3 Casimir-Lifshitz forces induced in RLC circuits: one in series, the other in parallel
As shown in the main text, the interaction force, $f^{s}_{RLC}$, induced by the RLC circuit in series, can be obtained from (S15)–(S19) by substituting $\Omega=\tfrac{1}{\sqrt{LC}}$, $\gamma=\tfrac{R}{L}$ and assuming that only the capacitance $C(d)$ depends on distance. As far as the RLC circuit in parallel is concerned, the interaction force $f^{p}_{RLC}$, induced by this circuit, is obtained from (S15)–(S19) by substituting $\Omega=\tfrac{1}{\sqrt{LC}}$, $\gamma=\tfrac{1}{RC}$, when only the inductance $L(d)$ depends on distance.
It follows from (S15) that the circuit-induced interaction force for the RLC circuit in series is
$f^{s}_{RLC}=\frac{T}{2C}\frac{\partial C}{\partial d}+\frac{i\hbar\frac{\partial C}{\partial d}}{2\pi C^{2}R\sqrt{\frac{4L}{CR^{2}}-1}}\left[\psi\left(1+\frac{\hbar R}{4\pi TL}\left(1-i\sqrt{\frac{4L}{CR^{2}}-1}\right)\right)-\right.\\ \left.-\psi\left(1+\frac{\hbar R}{4\pi TL}\left(1+i\sqrt{\frac{4L}{CR^{2}}-1}\right)\right)\right],$ (S20)
In the limit of weak dissipation $R\ll\sqrt{\tfrac{L}{C}},TL$ one gets
$f^{s}_{RLC}=\left[\frac{\hbar}{4\sqrt{LC^{3}}}\coth\frac{\hbar}{2T\sqrt{LC}}+\frac{\hbar^{2}R}{8\pi^{2}T(LC)^{\frac{3}{2}}}\operatorname{\mathrm{Im}}\left(\psi^{(1)}\biggl{(}1+\frac{i\hbar}{2\pi T\sqrt{LC}}\biggr{)}\right)\right]\frac{\partial C}{\partial d}.$ (S21)
At high temperatures the force is
$f^{s}_{RLC}=\left(\frac{T}{2C}+\frac{\hbar^{2}}{24TLC^{2}}\right)\frac{\partial C}{\partial d}.$ (S22)
In the low temperature limit we obtain
$f^{s}_{RLC}=\frac{\hbar\frac{dC}{dd}}{2\pi RC^{2}\sqrt{1-\frac{4L}{CR^{2}}}}\ln\frac{1+\sqrt{1-\frac{4L}{CR^{2}}}}{1-\sqrt{1-\frac{4L}{CR^{2}}}}.$ (S23)
Analogously, one gets for the RLC circuit in parallel
$f^{p}_{RLC}=\left\{\frac{T}{2L}+\frac{i\hbar R}{2\pi L^{2}\sqrt{\frac{4CR^{2}}{L}-1}}\left[\psi\left(1+\frac{\hbar}{4\pi TCR}\left(1-i\sqrt{\frac{4CR^{2}}{L}-1}\right)\right)-\right.\right.\\ \left.\left.-\psi\left(1+\frac{\hbar}{4\pi TCR}\left(1+i\sqrt{\frac{4CR^{2}}{L}-1}\right)\right)\right]\right\}\frac{\partial L}{\partial d},$ (S24)
In the limit of weak dissipation $R\gg\sqrt{\tfrac{L}{C}},\tfrac{1}{TC}$ one gets
$f^{p}_{RLC}=\left[\frac{\hbar}{4L^{\frac{3}{2}}C^{\frac{1}{2}}}\coth\frac{\hbar}{2T\sqrt{LC}}+\frac{\hbar^{2}}{8\pi^{2}T(LC)^{\frac{3}{2}}R}\operatorname{\mathrm{Im}}\left(\psi^{(1)}\biggl{(}1+\frac{i\hbar}{2\pi T\sqrt{LC}}\biggr{)}\right)\right]\frac{\partial L}{\partial d}.$ (S25)
At high temperatures the force is
$f^{p}_{RLC}=\left(\frac{T}{2L}+\frac{\hbar^{2}}{24TL^{2}C}\right)\frac{\partial L}{\partial d}.$ (S26)
In the low temperature limit we obtain
$f^{p}_{RLC}=\frac{\hbar R\frac{dL}{dd}}{2\pi L^{2}\sqrt{1-\frac{4CR^{2}}{L}}}\ln\frac{1+\sqrt{1-\frac{4CR^{2}}{L}}}{1-\sqrt{1-\frac{4CR^{2}}{L}}}.$ (S27)
S4 Casimir-like forces influenced by the frequency dispersion of the damping function
Within the Ohmic regime, the free energy expression (S8) diverges logarithmically at high frequencies. The same divergence also occurs in (1) for the interaction force when the damping function depends on the interbody distance but not on frequency. In such cases, one must take into account the frequency dispersion of the damping function. As in other similar cases, solvable specific models of dispersion demonstrate possible behaviors of the convergent results [51, 52, 53, 60].
Here, we use the Drude model $\gamma(\omega)=\frac{\gamma_{0}\omega_{D}}{\omega_{D}-i\omega}$, assuming that the three model parameters $\Omega$, $\gamma_{0}$, and $\omega_{D}$ may depend on the distance $d$ between the bodies.
In this case, the free energy (S8) takes the form
$F=T\ln\left[\dfrac{\hbar\Omega}{T}\prod_{n=1}^{\infty}\dfrac{(\omega_{n}+i\omega_{1})(\omega_{n}+i\omega_{2})(\omega_{n}+i\omega_{3})}{\omega_{n}^{2}(\omega_{n}+\omega_{D})}\right],$ (S28)
and, as follows from expression (1) in the main text of the paper, the interaction force is $f=f_{\Omega}+f_{\gamma}=f_{\Omega}+f_{\gamma_{0}}+f_{\omega_{D},1}+f_{\omega_{D},2}$, where
$f_{\Omega}=-2T\Omega\frac{\partial\Omega}{\partial d}\sideset{}{{}^{\prime}}{\sum}_{n=0}^{\infty}\frac{\omega_{n}+\omega_{D}}{(\omega_{n}+i\omega_{1})(\omega_{n}+i\omega_{2})(\omega_{n}+i\omega_{3})},$ (S29)
$f_{\gamma_{0}}=-T\frac{\partial\gamma_{0}}{\partial d}\sum_{n=1}^{\infty}\frac{\omega_{D}\omega_{n}}{(\omega_{n}+i\omega_{1})(\omega_{n}+i\omega_{2})(\omega_{n}+i\omega_{3})},$ (S30)
$f_{\omega_{D},1}=-T\frac{\partial\omega_{D}}{\partial d}\sum_{n=1}^{\infty}\frac{\gamma_{0}\omega_{n}}{(\omega_{n}+i\omega_{1})(\omega_{n}+i\omega_{2})(\omega_{n}+i\omega_{3})},$ (S31)
$f_{\omega_{D},2}=T\frac{\partial\omega_{D}}{\partial d}\times\\ \sum_{n=1}^{\infty}\frac{\omega_{D}\gamma_{0}\omega_{n}}{(\omega_{n}+\omega_{D})(\omega_{n}+i\omega_{1})(\omega_{n}+i\omega_{2})(\omega_{n}+i\omega_{3})}.$ (S32)
The complex eigenfrequencies of the oscillator $\omega_{1,2,3}$ appearing here satisfy the dispersion equation $\Omega^{2}-i\gamma(\omega)\omega-\omega^{2}=0$, which for the Drude model, takes the form
$\omega^{3}+i\omega_{D}\omega^{2}-(\Omega^{2}+\gamma_{0}\omega_{D})\omega-i\Omega^{2}\omega_{D}=0.$ (S33)
The roots $\omega_{1,2,3}$ of this equation must satisfy the relations
$\left\{\begin{aligned} &\omega_{1}+\omega_{2}+\omega_{3}=-i\omega_{D},\\ &\omega_{1}\omega_{2}+\omega_{1}\omega_{3}+\omega_{2}\omega_{3}=-(\Omega^{2}+\gamma_{0}\omega_{D}),\\ &\omega_{1}\omega_{2}\omega_{3}=i\Omega^{2}\omega_{D}.\end{aligned}\right$. (S34)
In this model, it is generally assumed that the Drude frequency is much larger than the other characteristic frequencies of the problem $\omega_{D}\gg\Omega,\gamma_{0}$. Retaining the zeroth-order ($\sim\omega_{D}$) and first-order ($\sim\Omega,\gamma_{0}$) terms yields an approximate solution [52]
$\left\{\begin{aligned} &i\omega_{1}=\frac{\gamma_{0}}{2}+i\sqrt{\Omega^{2}-\dfrac{\gamma_{0}^{2}}{4}},\\ &i\omega_{2}=\frac{\gamma_{0}}{2}-i\sqrt{\Omega^{2}-\dfrac{\gamma_{0}^{2}}{4}},\\ &i\omega_{3}=\omega_{D}-\gamma_{0}.\end{aligned}\right$. (S35)
From the third equation in (S34), it follows that one of the eigenfrequencies, hereafter denoted $\omega_{3}$, is of zero-order; i.e., its absolute value is comparable to $\omega_{D}$. The values $|\omega_{1,2}|$ are the first order terms, $\lesssim\max\left\{\Omega,\gamma_{0}\right\}\ll\omega_{D}$.
Solution (S35) satisfies the first relation in (S34) exactly, and the second and third relations approximately. Identifying the second-order term on the right-hand side of the second equation in (S34) requires knowledge of the second-order corrections to the second and third terms on its left-hand side: $\tilde{\omega}_{1}^{(2)}\tilde{\omega}_{3}^{(0)}+\tilde{\omega}_{2}^{(2)}\tilde{\omega}_{3}^{(0)}$. Therefore, the approximate solution (S35) reliably reproduces only the first-order term $\gamma_{0}\omega_{D}$ in this relation.
Regarding the third relation in (S34), solution (S35) reliably reproduces the second-order term on the right-hand side, since the product of all three solutions on the left-hand side of (S34) contains only zeroth- and first-order terms $\tilde{\omega}_{1}^{(1)}\tilde{\omega}_{2}^{(1)}\tilde{\omega}_{3}^{(0)}$ and does not include second-order corrections $\tilde{\omega}_{1,2}^{(2)}$. The solution (S35) satisfies the third relation in (S34) only approximately, leading to extra third-order terms, because the product of the first-order terms $\tilde{\omega}_{1}^{(1)}\tilde{\omega}_{2}^{(1)}\tilde{\omega}_{3}^{(1)}$ must be considered together with the products containing second-order corrections $\tilde{\omega}_{1}^{(2)}\tilde{\omega}_{2}^{(1)}\tilde{\omega}_{3}^{(0)}$ and $\tilde{\omega}_{1}^{(1)}\tilde{\omega}_{2}^{(2)}\tilde{\omega}_{3}^{(0)}$.
The validity of the first relation in (S34) ensures the expression of the infinite product over Matsubara frequencies in (S28) in terms of several Gamma functions (see, for example, [93], p. 7, Sec. 1.3, (8)):
$F$ $=-T\ln\left[\dfrac{T\Gamma\left(1+i\frac{\hbar\omega_{1}}{2\pi T}\right)\Gamma\left(1+i\frac{\hbar\omega_{2}}{2\pi T}\right)\Gamma\left(1+i\frac{\hbar\omega_{3}}{2\pi T}\right)}{\hbar\Omega\Gamma\left(1+\frac{\hbar\omega_{D}}{2\pi T}\right)}\right]$ (S36)
$=-T\ln\left[\hbar\Omega\dfrac{\Gamma\left(i\frac{\hbar\omega_{1}}{2\pi T}\right)\Gamma\left(i\frac{\hbar\omega_{2}}{2\pi T}\right)\Gamma\left(i\frac{\hbar\omega_{3}}{2\pi T}\right)}{4\pi^{2}T\Gamma\left(\frac{\hbar\omega_{D}}{2\pi T}\right)}\right].$
The second expression in (S36) follows from the first one, due to the equality $\Gamma(1+z)=z\Gamma(z)$ and the third relation in (S34).
The expression for the oscillator-induced Casimir-like force $f=-\left(\frac{\partial F}{\partial d}\right)_{T=\text{const}}$ can be found most easily by directly differentiating expression (S36). Alternatively, one can sum over Matsubara frequencies in (S29)-(S32) using known relations (see, for example, Ref. [95], p. 683, 5.1.24.6). Both methods lead to equivalent results within the accuracy of the approximation used. From the first expression in (S36), we find
$f=-\dfrac{T}{\Omega}\frac{\partial\Omega}{\partial d}-\frac{i\hbar}{2\pi}\left[\psi\biggl{(}1+i\frac{\hbar\omega_{2}}{2\pi T}\right$. $\left.\!\!\biggr{)}-\psi\left(1+i\frac{\hbar\omega_{1}}{2\pi T}\right)\right]\frac{\left(\Omega\dfrac{\partial\Omega}{\partial d}-\dfrac{\gamma_{0}}{4}\dfrac{\partial\gamma_{0}}{\partial d}\right)}{\sqrt{\Omega^{2}-\frac{\gamma_{0}^{2}}{4}}}+\!\!$ (S37)
$+\dfrac{\hbar}{4\pi}\left[\psi\left(1+i\frac{\hbar\omega_{1}}{2\pi T}\right)+\right$. $\left.\psi\left(1+i\frac{\hbar\omega_{2}}{2\pi T}\right)-2\psi\left(1+\frac{\hbar(\omega_{D}-\gamma_{0})}{2\pi T}\right)\right]\dfrac{\partial\gamma_{0}}{\partial d}+$
$+\dfrac{\hbar}{2\pi}$ $\left[\psi\left(1+\frac{\hbar(\omega_{D}-\gamma_{0})}{2\pi T}\right)-\psi\left(1+\frac{\hbar\omega_{D}}{2\pi T}\right)\right]\dfrac{\partial\omega_{D}}{\partial d}.$
Formulas (4) and (5) in the main text describe the same result after replacing $\gamma$ with $\gamma_{0}$ in (3).
Using the expressions (S36) and (S37), the free energy and Casimir-like force can be easily described in the limiting cases of high and low temperatures, as well as for different possible relationships between temperature and Drude frequency. Here we present the corresponding results for the Casimir-like force.
For a very high temperature $\Omega,\gamma_{0}\ll\omega_{D}\ll T$, when the parameter $\frac{\omega_{D}}{T}$ is small, one obtains from (S37) the following dominant terms in each of the three groups of terms, after expanding the digamma function near unity:
$f=-\dfrac{T}{\Omega}\dfrac{\partial\Omega}{\partial d}-\dfrac{\hbar^{2}\omega_{D}}{24T}\dfrac{\partial\gamma}{\partial d}-\dfrac{\hbar^{2}\gamma}{24T}\dfrac{\partial\omega_{D}}{\partial d}$ (S38)
If the temperature is high only in comparison to $\Omega$ and $\gamma$, while $\omega_{D}\gg 2\pi T\gg\Omega,\gamma$, then, in the presence of a large value of the argument $\frac{\omega_{D}}{2\pi T}$, the asymptotic expansion $\psi(1+z)=\ln z+\frac{1}{2z}+\ldots$ should be used. For the oscillator-induced Casimir-like force, this yields expression (6) in the main text.
At low temperatures, the arguments of all functions in (S37) take large values. Using the corresponding asymptotic expansions and keeping the dominant terms in each of the three groups of terms containing derivatives $\frac{\partial\Omega}{\partial d}$, $\frac{\partial\gamma_{0}}{\partial d}$ and $\frac{\partial\omega_{D}}{\partial d}$, we arrive at the following expression:
$f=-\frac{\hbar\Omega}{2\pi\sqrt{\frac{\gamma_{0}^{2}}{4}-\Omega^{2}}}\ln\frac{\frac{\gamma_{0}}{2}+\sqrt{\frac{\gamma_{0}^{2}}{4}-\Omega^{2}}}{\frac{\gamma_{0}}{2}-\sqrt{\frac{\gamma_{0}^{2}}{4}-\Omega^{2}}}\frac{\partial\Omega}{\partial d}-\Biggl{[}\frac{\hbar}{2\pi}\ln\frac{\omega_{D}}{\Omega}-\frac{\hbar\gamma_{0}}{8\pi\sqrt{\frac{\gamma_{0}^{2}}{4}-\Omega^{2}}}\ln\frac{\frac{\gamma_{0}}{2}+\sqrt{\frac{\gamma_{0}^{2}}{4}-\Omega^{2}}}{\frac{\gamma_{0}}{2}-\sqrt{\frac{\gamma_{0}^{2}}{4}-\Omega^{2}}}\Biggr{]}\frac{\partial\gamma_{0}}{\partial d}-\,\frac{\hbar\gamma_{0}}{2\pi\omega_{D}}\frac{\partial\omega_{D}}{\partial d}.$ (S39)
Under the condition $2\Omega>\gamma_{0}$, it is convenient to transform (S39) into the following form
$f=-\frac{\hbar\Omega}{2\sqrt{\Omega^{2}-\frac{\gamma_{0}^{2}}{4}}}\left(1-\frac{2}{\pi}\arctan\frac{\gamma_{0}}{2\sqrt{\Omega^{2}-\frac{\gamma_{0}^{2}}{4}}}\right)\frac{\partial\Omega}{\partial d}-\\ -\left[\frac{\hbar}{2\pi}\ln\frac{\omega_{D}}{\Omega}-\frac{\hbar\gamma_{0}}{8\sqrt{\Omega^{2}-\frac{\gamma_{0}^{2}}{4}}}\left(1-\frac{2}{\pi}\arctan\frac{\gamma_{0}}{2\sqrt{\Omega^{2}-\frac{\gamma_{0}^{2}}{4}}}\right)\right]\frac{\partial\gamma_{0}}{\partial d}-\frac{\hbar\gamma_{0}}{2\pi\omega_{D}}\frac{\partial\omega_{D}}{\partial d}.$ (S40)
It can be verified that the factor preceding the term $\frac{\partial\gamma_{0}}{\partial d}$ in (S39) is positive throughout the domain of applicability of the derived expression $\omega_{D}\gg\Omega,\gamma$. Therefore, in accordance with the result stated in the main text, each of the three contributions to the Casimir force in (S39) describes attraction, when the corresponding parameter $\Omega$, $\gamma_{0}$, or $\omega_{D}$ increases with $d$, and repulsion when it decreases with $d$.